OpenStudy (anonymous):

Evaluate x^-2 for x = -3. I got 1/9 but apparently it was wrong

jimthompson5910 (jim_thompson5910):

\[\Large x^{-2} = \frac{1}{x^2}\] \[\Large (-3)^{-2} = \frac{1}{(-3)^2}\] \[\Large (-3)^{-2} = \frac{1}{9}\] I'm getting the same

OpenStudy (anonymous):

i'll go with that too

OpenStudy (anonymous):

@jim_thompson5910 @satellite73 Thanks for trying to help (:

jimthompson5910 (jim_thompson5910):

It sounds like there's a typo somewhere. So I'd just ask the teacher and s/he will probably agree and give points back.

OpenStudy (misty1212):

me too, i would go with \(\frac{1}{9}\)

OpenStudy (misty1212):

in fact, i am sure of it

OpenStudy (triciaal):

do they want it expressed as 9^-1?

OpenStudy (loser66):

If every body failed on this problem, I want to join to the group. :)